### v3

# Supplementary Material For Self-supervised Likelihood Estimation with Energy Guidance for Anomaly Segmentation in Urban Scenes

Supplementary Material For Self-supervised Likelihood Estimation with Energy Guidance for Anomaly Segmentation in Urban Scenes
Anonymous submission
More Experimental Results
Architecture Method FS LAF Val FS Static Val mIoU
$\rm{FPR_{95}}$downarrow AP $\uparrow$ $\rm{FPR_{95}}$downarrow AP $\uparrow$
OCRNet (YuanCW20) SML 18.28 39.96 15.07 47.90 80.66
JEM 22.90 23.27 16.80 34.03
Void Class 16.65 46.62 17.74 29.30
SLEEG 8.89 72.51 7.6 73.01
ISANet (huang2019isa) SML 18.67 28.76 14.86 32.15 80.51
JEM 35.57 22.65 16.22 45.22
Void Class 29.17 43.53 17.18 24.61
SLEEG 12.28 65.79 3.33 80.03
FCN (shelhamer2017fully) SML 39.80 17.59 14.53 28.92 77.72
JEM 39.36 21.83 13.47 30.51
Void Class 24.29 43.44 14.79 22.72
SLEEG 21.01 63.11 5.31 61.29
Table 1: Comparison between SLEEG and other anomaly detection methods on FS validation sets with different segmentation models. ”Void Class” denotes training models with pixels that fall into void class as anomaly samples.
Figure 1: Visualization of on FS Lost & Found validation set. Compared with JEM, predictions from our SLEEG show anomaly maps with higher responses for anomalous instances and lower values for normal pixels.
Results on Lost & Found. To make more comprehensive comparison, we also evaluate our method on the old Lost & Found dataset (lostAndFound2016), where 13 challenging real-world scenes are included with 37 different anomalous instances, which is the first publicly available urban anomaly segmentation datasets. Specifically, the obstacles in this dataset vary significantly in size and material. And following the official protocol (lostAndFound2016), 1,203 images collected from 112 video stereo sequences with a resolution of 2048$\times$1024 are utilized as the test set. Similar to the setting on FS Lost & Found, the segmentation model of DeepLab series (chen2018encoder) with ResNet101 pretrained on Cityscapes is used and fixed without re-training on this dataset. Finally, the same OoD head and evaluation metrics are adopted as FS Lost & Found.
[…]
Meta-OoD (chan2021entropy) 97.95 71.23 5.95
${}^{\dagger}$ (di2021pixel) 98.38 70.43 4.89
Deep ${}^{\dagger}$ (liu2019deep) 98.67 72.73 3.81
${}^{\dagger}$ (di2021pixel) 99.76 78.29 0.81
${}^{\dagger}$ (grcic2022densehybrid) 99.37 78.70 2.10
MSP (hendrycks2019scaling) 85.49 38.20 18.56
Mahalanobis (lee2018simple) 79.53 42.56 24.51
Max Logit (hendrycks2016baseline) 94.52 65.45 15.56
Entropy (hendrycks2016baseline) 86.52 50.66 16.95
${}^{\dagger}$ (jung2021standardized) 88.05 25.89 44.48
Energy (liu2020energy) 94.45 66.37 15.69
SLEEG 98.59 82.88 4.60
Table 2: Comparison on Lost & Found testing set. All methods use the same segmentation models. $\dagger$ indicates training with WideResNet38 backbone. Bold values and underlined values represent the best and second best results.
Tab. 2 shows the performance of the proposed SLEEG on the test set of Lost & Found. Notably, our method surpasses all previous approaches and achieves SOTA performance on AP. Specifically, though previous SOTA methods DenseHybrid (grcic2022densehybrid) and PEBAL (tian2021pixel) utilize auxiliary data and are re-trained with more complex WideResNet38 as backbone, our SLEEG still brings a relative improvement of 4.59% and 4.18% to them on AP respectively with ResNet101 as backbone and requires no auxiliary OoD samples or further re-training of the segmentation model. Furthermore, when compared with the SOTA method Energy (liu2020energy) that falls into the same category as SLEEG, our method surpasses it by a large gap of 16.51% on AP and meanwhile significantly reduces the false positive pixels from 15.69% to 4.6%. This results fully illustrate that SLEEG shows great robustness and effectiveness on localising various real-world unexpected objects. Moreover, SML (jung2021standardized) attempts to tackle this task by re-balance the class-wise discrepancy of inlier samples. However, since there are only two classes in Lost & Found, there exists large performance gap between the performance of SML on Fishyscapes and Lost & Found, implying that SML may not be suitable for real-world applications. By contrast, it is worth noting that our SLEEG is capable of achieving consistent performance boost on all these datasets, demonstrating its great potential in real world scenarios.
Method FS LAF Validation
$\rm{FPR_{95}}$downarrow AP $\uparrow$ AUROC $\uparrow$
DenseHybrid + OoD data  (grcic2022densehybrid) 6.1 63.8 -
DenseHybrid + Convex patch 22.2 56.2 96.6
DenseHybrid + Refined convex patch 18.6 60.5 97.0
SLEEG + Refined convex patch 10.9 70.9 98.3
Table 3: Investigation when combining other state-of-the-art method DenseHybrid (grcic2022densehybrid) with our strategy of patch generation on FS LAF validation set.
Table 4: Performance evaluation on the SMIYC benchmark (chan21arxiv).
Method OoD Re- AnomalyTrack
Data training $\mathrm{FPR}_{95}$downarrow AP $\uparrow$
SynBoost (biase21cvpr) 61.9 56.4
JSRNet (vojir21iccv) 43.9 33.6
Void Classifier (blum21ijcv) 63.5 36.6
DenseHybrid (grcic2022densehybrid) 9.8 78.0
PEBAL (tian2021pixel) 40.8 49.1
Image Resyn. (lis19iccv) 25.9 52.3
Embed. Dens. (blum21ijcv) 70.8 37.5
ODIN (liang18iclr) 71.7 33.1
MC Dropout (kendall17nips) 69.5 28.9
Max softmax (hendrycks17iclr) 72.1 28.0
Mahalanobis (lee18nips) 87.0 20.0
SLEEG (ours) 36.4 52.7
Table 5: Performance evaluation on StreetHazards (streethazards2019). SLEEG achieves competitive anomaly detection performance.
Method OoD Re- Anomaly detection
Data training $\mathrm{FPR}_{95}$downarrow AUROC $\uparrow$ AP $\uparrow$
Energy (liu20neurips) 18.2 93.0 12.9
Outlier Exposure (hendrycks19iclr) 17.7 94.0 14.6
OOD-Head (bevandic19gcpr) 56.2 88.8 19.7
OH-MSP (bevandic21arxiv) 30.9 89.7 18.8
DenseHybrid (grcic2022densehybrid) 13.0 95.6 30.2
CoroCL (liu2023residual) 8.2 97.2 31.2
SynthCP (xia20eccv) 28.4 88.5 9.3
Dropout (kendall17nips)(xia20eccv) 79.4 69.9 7.5
TRADI (franchi20eccv) 25.3 89.2 7.2
OVNNI (franchi20arxiv) 22.2 91.2 12.6
SO+H (grcic21visapp) 25.2 91.7 12.7
DML (cen21iccv) 17.3 93.7 14.7
MSP (hendrycks17iclr) 27.9 90.1 7.5
ML (hendrycks19arxiv) 22.5 92.4 11.6
ODIN (liang18iclr) 28.7 90.0 7.0
ReAct (sun21neurips) 21.2 92.3 10.9
SLEEG (ours) 13.3 95.7 27.6
Figure 2: Investigation on the influence on AP and false positive rate with varied $\lambda$ value on FS Lost & Found validation set (left) and FS Static validation set (right).
Results for OoD data replacement. Meanwhile, we also separately investigate our patch policy with other method (grcic2022densehybrid) by replacing the external OoD data used in original paper with our generated patches. The results can be illustrated in Tab. 3, where the results of Densehybrid+OoD data is borrowed from the original paper, and the other results are reimplemented by ourselves. It can be observed that although applying less data, our patch policy can achieve results close to those in the original paper, this further demonstrate that our scheme can be viewed as a potential replacement for OoD data relied by previous methods.
Parameter Sensitivity. Finally we investigate the influence of the balance factor $\lambda$ on the FS Lost & Found validation set and FS Static validation set. As shown in Fig. 2, SLEEG performs best when the balance factor $\lambda=0.5$. The results demonstrate that SLEEG achieves similar performance across all settings, implying SLEEG is not very sensitive to $\lambda$. Finally, we set $\lambda=0.5$.
Figure 3: Visualization results of some failure cases from SLEEG on FS Static validation set.
Qualitative Results
We also conduct further comparison with the SOTA method DenseHybrid on Lost & Found test set and FS Lost & Found validation set. As shown in Fig. 4, for the challenging small anomalous objects at a distance, SLEEG produces larger anomaly scores, allowing vehicles respond to prevent accidents in a more timely manner. Besides, for objects with diverse shapes, our SLEEG is capable of producing a more complete contour covering the entire anomalous object as shown in the third row of Fig. 4.
Fig. 6 further visualizes the comparison between JEM and our SLEEG on Lost & Found test set. Notably, SLEEG can consistently perform better than JEM when detecting objects with various scales and distances. Beside JEM, Additionally, Fig. 6 and Fig. 5 describe the comparison of visualization results between the classic softmax entropy (MSP) and image re-synthesis (imgr) methods and our SLEEG. It can be observed that these methods perform much worse than our SLEEG and are unable to effectively distinguish in-distribution objects from anomalous ones. And there even exists large background regions that are assigned with higher values than anomalous objects. By contrast, our SLEEG generates anomaly maps where unexpected instances are intensively assigned with higher scores, resulting in much fewer false positive pixels as well.
Limitations
As shown in the Table 1 of the manuscript and Fig. 3, our SLEEG performs inferior to SOTA methods that utilize auxiliary OoD data on the FS Static dataset. This mainly accounts for that the synthetic anomalous objects in FS Static are similar to the instances in the auxiliary dataset that these methods use, i.e., COCO (lin2014microsoft) and ADE20K (zhou2019semantic). Therefore, training with such OoD samples allows models foresee the anomalous objects, resulting in their good performance. Moreover, training in this manner also results in a large gap between these methods and ours in performance on the FS Lost & Found and Lost & Found datasets since anomalous objects in the FS Lost & Found dataset are more realistic. Finally, though our SLEEG performs inferior on FS Static dataset, since the anomaly segmentation task targets at real-world automatic driving scenarios, the ability of tackling realistic unexpected samples is more essential for our SLEEG. Therefore, we focus on utilizing the inherent spatial context to improve the ability of tackling with the distribution discrepancy of normal and anomalies without requiring extra OoD data.
Figure 4: Comparison of visualization results with Densehybrid on Lost & Found test set and FS Lost & Found validation set. The proposed SLEEG is capable of localizing small anomalous objects more accurately. For example, SLEEG generates higher anomaly scores for the instances in the red box of the second and third rows.
Figure 5: Comparison of visualization results with JEM, Softmax Entropy and Image Re-synthesis on FS Lost & Found validation set.
Figure 6: Comparison of visualization results with JEM, Softmax Entropy and Image Re-synthesis on Lost & Found test set.